# ObfuscaTune: Obfuscated Offsite Fine-tuning and Inference of Proprietary LLMs on Private Datasets

### v1

@@ 4 Conclusion
This work tackled the timely but underexplored problem of performing offsite inference and finetuning of a proprietary LLM owned by a model provider entity on the confidential/private data of another data owner entity, in a way that ensures the confidentiality of both the model and the data. Our proposed approach, ObfuscaTune, achieves this by combining a simple yet effective obfuscation technique with an efficient usage of confidential computing (only ${}5\%$ of the model parameters are placed on TEE). Our extensive empirical evaluation on four NLP benchmark datasets and different models highlights the effectiveness of our method and emphasizes the importance of using random matrices with low condition numbers for preserving high utility. In future work, we will investigate the effectiveness of our approach to RAG-systems.
5 Limitations
One potential limitation of our work is that despite testing on different models and datasets, we focused on the same model architecture, i.e., GPT2. However, most of the other LLMs are composed on the same building blocks, which makes the application of our method to them straightforward. Another limitation might be that while the slowdown incured by ObfuscaTune is substantially lower than other technologies, e.g., MPC and HE, it might still be unsuitable for some applications where efficiency has a higher importance than privacy
@@ Appendix A Hyperparameters
We train all models for 10 epochs. We perform validation at the end of every epoch and use early stopping with a patience of 3. We use a learning rate of ${3e-5}$ and a batch size of 1. We keep the other hyperparameters unchanged from Karpathy (2023). For LoRA, we use the hyperparameters: $r=16$, $\alpha=32$ and apply dropout with $0.05$. We did not perform hyperparameter tuning, which highlights the robustness of our method. We did all experiments on middle-range GPUs. Each experiment took between less than 1 and 8 GPU hours, depending on he model size and dataset.
Figure 1: Overview of the proposed ObfuscaTune, composed by the three stakeholders: model provider, which seeks to keep the model confidential, data owner, which uses the model (finetuning or inference) while preserving privacy of their data, and cloud provider which provides the computation infrastructure, while potentially trying to eavesdrop on the data or steal the model. ObfuscaTune provides the necessary protection by keeping very few components of the model within a TEE, and obfuscating the remaining ones, effectively and efficiently preventing data or model stealing. This Figure will be part of the additional page in the camera ready version upon paper acceptance.
Figure 2: Detailed architecture of the GPT-2 with M layers using ObfuscaTune. Diagram blocks in green are within the TEE, while the orange are outside the TEE. This diagram illustrates how the data is successfully sent from and to the TEE, while being obfuscated while outside the TEE. Note that both the input text and output text are always within the TEE to prevent inversion attacks. Note that the non-activation applied after the first MLP (bottom) is applied on the de-obfuscated embedding. The same applies for the softmax non-linear function.
Appendix B Effect of the condition number
The condition number $\kappa$ of a matrix $A$ is defined as $\kappa(A)=\frac{M}{m}$, where $M=\max{\frac{\|Ax\|}{\|x\|}}$ measures how much the mapping induced by that matrix can stretch vectors and $m=\min{\frac{\|Ax\|}{\|x\|}}$ measures how much it can shrink vectors. It determines how much a relative error in the input reflects on the output for solving linear systems, matrix inversion or matrix-vector multiplication Golub and Van Loan (2013). Such numerical errors get accumulated and increase with the number of sequential matrix multiplication operations, i.e., the deeper the model the higher the accumulated error. We minimize the numerical errors by minimizing the condition number of the random matrix.
In this work, we consider the condition number w.r.t the $\ell_{2}$ norm. Since orthogonal matrices induce isometries, i.e $\|Ax\|_{2}=\|x\|_{2}$ for all $x$, we get $\kappa(A)=1$ for every orthogonal matrix $A$. Note that singular matrices have the highest (worst) possible condition number, which is $\infty$, since for a singular matrix $A$, $m=\min{\frac{\|Ax\|}{\|x\|}}=0$. On the other side, from the definition we see that the lowest possible $\kappa$ is 1.
Let $\sigma_{max}(A)$ and $\sigma_{min}(A)$ respectively be the largest and the smallest singular values of the matrix $A$. For the $\ell_{2}$-induced operator norm norm the following holds :
@@ Appendix B Effect of the condition number
The last equation makes it possible to generate random matrices $R$ of a given predefined condition number $\kappa(R)$. First we generate random matrices $A$ and $B$ using the standard normal distribution. We then apply QR-decomposition on $A$ and $B$ to generate two orthogonal matrices $Q_{A}$ and $Q_{B}$. We then choose a random positive value for the largest singular value of the final matrix $R$ and we set $\sigma_{min}(R)=\frac{\sigma_{max}(R)}{\kappa(R)}$. The remaining singular values can be sampled randomly from the uniform distribution between $\sigma_{min}(R)$ and $\sigma_{max}(R)$. Then we construct the diagonal matrix $S$ with the singular values on the diagonal. Note that $S^{-1}$ is the diagonal matrix with the inverses of the singular values on the diagonal. Then we define $R$ to be having the following singular value decomposition:
@@
And can calculate $R^{-1}=Q_{B}^{T}S^{-1}Q_{A}^{T}$ with minimal rounding errors. We use this approach to generate random matrices of a given condition number and monitor the effect of the condition number on the test accuracy of the final model. The results are showcased in table 2 show indeed that it is curcial to have a low condition number, otherwise the training degenerates.

### v2

@@ 6 Conclusion
This work tackled the timely but underexplored problem of performing offsite inference and finetuning of a proprietary LLM owned by a model provider entity on the confidential/private data of another data owner entity, in a way that ensures the confidentiality of both the model and the data. Our proposed approach, ObfuscaTune, achieves this by combining a simple yet effective obfuscation technique with efficient use of confidential computing (only $\leavevmode\nobreak\ 5\%$ of the model parameters are placed on TEE). Our extensive empirical evaluation on four NLP benchmark datasets and different models highlights the effectiveness of our method and emphasizes the importance of using random matrices with low condition numbers to preserve high utility. In future work, we will investigate the effectiveness of our approach to RAG-systems.
@@ Appendix A Hyperparameters
We train all models for 10 epochs. We perform validation at the end of every epoch and use early stopping with a patience of 3. We use a learning rate of ${3e-5}$ and a batch size of 1. We keep the other hyperparameters unchanged from [26]. For LoRA, we use the hyperparameters: $r=16$, $\alpha=32$ and apply dropout with $0.05$. We did not perform hyperparameter tuning, which highlights the robustness of our method. We did all experiments on middle-range GPUs. Each experiment took between 1 and 8 GPU hours, depending on the model size and dataset.
Appendix B Random matrix generation with a pre-defined condition number
Let $\sigma_{max}(A)$ and $\sigma_{min}(A)$, respectively, be the largest and the smallest singular values of the matrix $A$. For the $\ell_{2}$-induced operator norm norm the following holds :
@@ Appendix B Random matrix generation with a pre-defined condition number
To generate random matrices $R$ of a given predefined condition number $\kappa(R)$, we leverage this last equation as follows. First, we generate random matrices $A$ and $B$ using the standard normal distribution. We then apply QR-decomposition on $A$ and $B$ to generate two orthogonal matrices $Q_{A}$ and $Q_{B}$. We then choose a random positive value for the largest singular value of the final matrix $R$ and set $\sigma_{min}(R)=\frac{\sigma_{max}(R)}{\kappa(R)}$. The remaining singular values can be sampled randomly from the uniform distribution between $\sigma_{min}(R)$ and $\sigma_{max}(R)$. Then we construct the diagonal matrix $S$ with the singular values on the diagonal. Note that $S^{-1}$ is the diagonal matrix with the inverses of the singular values on the diagonal. Then we define $R$ to be having the following singular value decomposition:
@@
And can calculate $R^{-1}=Q_{B}^{T}S^{-1}Q_{A}^{T}$ with minimal rounding errors. We use this approach to generate random matrices of a given condition number and monitor the effect of the condition number on the test accuracy of the final model. The results presented in table 2 show indeed that it is crucial to have a low condition number, otherwise the training degenerates.